An experiment shows that a 114 mL gas sample has a mass of 0.172 g at a pressure of 727 mmHg and a temperature of 32 ∘C.
1
Expert's answer
2018-03-19T08:52:04-0400
Answer:
T = 32 + 273 = 305 K
p = 729/760 = 0.956 mmHg
V = 0.114 litre
n = pV/RT
n = 0.956×0.114/0.08206×305 = 0.0435 mole
M = m/n
M = 0.172/0.0435 = 39.5 g/mole